OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, –3) and has a slope of 2? Answer 2x – y = –5 2x + y = 5 2x + y = –5 2x – y = 5

OpenStudy (anonymous):

2x-y=5

OpenStudy (anonymous):

(1, –3) and has a slope of 2 y + 3 = 2 (x -1) y + 3 = 2x - 2 -> 2x -3 = 5

OpenStudy (anonymous):

Oop 2x - y = 5